OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the points (-1,-4) and (1,4)? write the equation in slope-intercept form

OpenStudy (ranga):

Do you know the formula for the slope of a line connecting two points?

OpenStudy (anonymous):

no that's all it said

OpenStudy (ranga):

slope m = (y2 - y1) / (x2 - x1) The two points are: (-1,-4) and (1,4) so x1 = -1, y1 = -4, x2 = 1, y2 = 4 Just plug the values into the formula and let me know what you get for m.

OpenStudy (anonymous):

I got 8/-2 is that right?

OpenStudy (anonymous):

can you help?

OpenStudy (ranga):

The slope m = ( 4 - (-4) ) / ( 1 - (-1) ) = 8 / 2 = 4 The equation of a line is: y = mx + b = 4x + b We know the point (1, 4) lies on it; put x = 1 and y = 4 and solve for b 4 = 4(1) + b. So b = 0 The equation of the line is: y = 4x